OpenStudy (anonymous):

compute the length of the curve from 0<= x <=3 of the function y = 1/3(x^2 +2)^3/2

OpenStudy (anonymous):

just grab a measuring tape :DD

ganeshie8 (ganeshie8):

arc length = \(\large \int_0^3 \sqrt{1 + (\frac{dy}{dx})^2} dx \)

ganeshie8 (ganeshie8):

find the derivative and plugin

ganeshie8 (ganeshie8):

\(\large y = \frac{1}{3}(x^2 +2)^{\frac{3}{2}}\) \(\large \frac{dy}{dx} = ?\)

OpenStudy (anonymous):

ok... here goes! sorry, I skipped it and got caught up in another.

ganeshie8 (ganeshie8):

yah

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= (\frac{ 1 }{ 3 })(\frac{ 3 }{ 2 })(x^2+2)^{1/2}(2x) = \frac{ 1 }{ 2}(2x)(x^2+2)^{1/2}\]

OpenStudy (anonymous):

well also the 1/2 cancels with the 2x

zepdrix (zepdrix):

Looks good bro. Plug it all in. Jam through it.